# Extremal simplicial distributions on cycle scenarios with arbitrary outcomes

### v1

@@ 1 Introduction
Simplicial distributions [1] are a new framework for studying contextuality that extends the sheaf-theoretic framework of Abramsky–Brandenburger [2] and the cohomological framework of Okay et al. [3]. Measurements and outcomes are represented by combinatorial models of spaces known as simplicial sets, and the non-signaling conditions are encoded topologically. The theory of simplicial distributions is an extension of the theory of non-signaling distributions, which utilizes topological ideas and tools for studying the polytopes of non-signaling distributions and the Bell polytopes describing the classical region [4]. In this paper, we use homotopical techniques to identify the extremal distributions on cycle scenarios with arbitrary outcomes that extend earlier results in the literature. Our techniques generalize to more sophisticated measurement spaces obtained by gluing cycle scenarios.
@@ 2.1 Simplicial distributions
  • An $n$-circle is a simplicial set $C^{(n)}$ specified by a sequence of pairwise distinct $1$-simplices $\sigma_{1},\cdots,\sigma_{n}$ satisfying $d_{0}(\sigma_{i})=d_{1}(\sigma_{i+1})$ for every $11\leq i\leq n-$ 1 and $d_{0}(\sigma_{n})=d_{1}(\sigma_{1})$. We will write $v_{i}=d_{1}(\sigma_{i})$.
@@
Figure 1: The $4$-circle measurement space.
More generally, we will consider $1$-dimensional measurement spaces, i.e., basically directed graphs. In this case the simplicial set representing the measurement space is specified by its 00-simplices and $1$-simplices together with the face maps between them. We will write a distribution on an edge $\sigma$ as a $d\times d$-matrix $M$ where $M_{ab}=p^{ab}$. For convenience we index the rows and columns of such a matrix by $10,\cdots,d-$ 1. The marginals on the initial vertex and the terminal vertex are given by the sum of the rows and columns, respectively.
@@ 2.2 The vertex support
To prove (1) suppose that $\varphi\in\operatorname{supp}(p)$, i.e., $\delta^{\varphi}\preceq p$. By Corollary 2.6 this implies that $p=\delta^{\varphi}$, a contradiction since $p$ is contextual. Part (2) is proved in [10, Pro. 2.12]. ∎
@@ 2.3 Null-homotopy
We have a simplicial map $d_{0}:\text{Dec}^{0}(Y)\to Y$ that sends a simplex to its 00-th face. For $Y$ we will take the nerve space $N\mathbb{Z}_{d}$. The set of $n$-simplices of the nerve space consists of $n$-tuples of elements $(a_{1},a_{2},\cdots,a_{n})\in\mathbb{Z}_{d}^{n}$ together with the face and the degeneracy maps:
$d_{i}(a_{1},\cdots,a_{n})$ =$=\left\{\begin{array}[]{ll}(a_{2},\cdots,a_{n})&i=0\\ (a_{1},\cdots,a_{i}+a_{i+1},\cdots,a_{n})&0<i<n\\ (a_{1},\cdots,a_{n-1})&i=n\end{array}\right.$
@@ 2.4 Vertices of faces
${D(\Delta_{\mathbb{Z}_{d}})}$${X}$${D(N\mathbb{Z}_{d})}$$\scriptstyle{D(\kappa)}$$\scriptstyle{p}$$\scriptstyle{\delta^{\varphi}}$
@@ 3.2 Morphisms of bundle scenarios
[…]
@@
Part (2): Suppose that $p\in\operatorname{sDist}(f)$ is a vertex. If there exist $q,s\in\operatorname{sDist}(g)$ and $10<\alpha<$ 1 such that $i_{\ast}(p)=\alpha q+(1-\alpha)s$, then by Lemma 3.4 we have
@@
$i_{\ast}(p)=\alpha i_{\ast}(\tilde{q})+(1-\alpha)i_{\ast}(\tilde{s})=i_{\ast}(\alpha\tilde{q}+(1-\alpha)\tilde{s}).$
@@ 4 Cycle scenario with arbitrary outcome
According to Proposition 2.11 the map $\varphi$ is not null-homotopic. Let $p$ be a distribution in $\operatorname{Face}(\varphi)$, i.e., $D(\kappa)(p)=\delta^{\varphi}$. For $11\leq i\leq n-$ 1 we have $D(\kappa)(p)_{\sigma_{i}}=\delta^{0}$, i.e., $\sum_{a=0}^{k-1}p_{\sigma_{i}}^{aa}=1$, which means that $p_{\sigma_{i}}^{ab}=0$ if $b\neq a$. In addition, $D(\kappa)(p)_{\sigma_{n}}=\delta^{1}$, and thus $\sum_{a=0}^{k-1}p_{\sigma_{n}}^{a(a+1)}=1$, that means $p_{\sigma_{n}}^{ab}=0$ if $b\neq a+1$. So
@@
and $p_{\sigma_{n}}^{(k-1)0}=p_{d_{0}(\sigma_{n})}^{0}=p_{d_{1}(\sigma_{1})}^{0}=p_{\sigma_{1}}^{00}$. We conclude that $p_{\sigma_{i}}^{aa}=\frac{1}{k}$ for every $11\leq i\leq n-$ 1 and $p_{\sigma_{n}}^{a(a+1)}=\frac{1}{k}$. We proved that the only distribution in $\operatorname{Face}(\varphi)$ is the $k$-cycle distribution that corresponds to the following sequence
@@
Corollary 4.6 generalizes Theorem 1 from [8]. In that reference the authors describe the contextual (non-local) extremal distributions (vertices) on the $4$-circle. Our result above extends this characterization to $n$-circle spaces with arbitrary outcomes.
@@ 5.1 Popescu-Rohrlich boxes
  • $A$ consists of a single edge ($1$-simplex) $\sigma_{1}$, and
@@
Figure 2: Decomposition of the $4$-circle.
Since $A$ is a $1$-simplex and $B$ is a directed path obtained by gluing along 00-simplices it is known (see, e.g., [1, Example 3.11, Corollary 4.6]) that any distribution on $(A,\Delta_{\mathbb{Z}_{d}})$ and $(B,\Delta_{\mathbb{Z}_{d}})$ is noncontextual, thus the vertices are all deterministic. It follows that $\operatorname{Vsupp}(p|_{K}){\cong}\operatorname{supp}(p|_{K})$ for $K=A,B$. So one can see that
@@ 5.2 A scenario with trichotomic measurements
Figure 3: The measurement space decomposed into two parts given by $2$-circles.
@@ 5.3 $(2,3,3)$ Bell scenario
[…]
@@ 6 Conclusion
We solve the vertex enumeration problem for the polytope of simplicial distributions on cycle scenarios with arbitrary outcomes. The simplicial distribution formulation aligns with the usual non-signaling formulation. Our approach uniquely benefits from novel homotopical methods, extending to scenarios formed by gluing cycle scenarios together. This gluing process is a crucial aspect of the theory of simplicial distributions when studying contextual distributions.

### v2

@@ 1 Introduction
Simplicial distributions [1] are a recent framework for studying contextuality that extends the sheaf-theoretic framework of Abramsky–Brandenburger [2] and the cohomological framework of Okay et al. [3]. Measurements and outcomes are represented by combinatorial models of spaces known as simplicial sets, and the non-signaling conditions are encoded topologically. The theory of simplicial distributions is an extension of the theory of non-signaling distributions, which utilizes topological ideas and tools for studying the polytopes of non-signaling distributions and the Bell polytopes describing the classical region [4]. In this paper, we apply homotopical techniques to fully characterize the extremal distributions on cycle scenarios with arbitrary outcomes, thereby extending the well-known binary-outcome case. Our techniques generalize to more sophisticated measurement spaces obtained by gluing cycle scenarios.
@@
This result provides a complete classification of extremal simplicial distributions on cycle scenarios with arbitrary outcome sets. The contextual extremal points correspond precisely to $k$-order cycle distributions with $k\geq 2$, while the remaining extremal points are deterministic and correspond to the case $k=1$. The key novelty lies in allowing arbitrary outcome sizes: this extends the well-studied binary-outcome case and yields a full solution to the extremal problem for cycle scenarios with arbitrary outcomes.
Leveraging the explicit form of $k$-order cycle distributions, we also derive a formula for the total number of vertices of the polytope of simplicial distributions associated with the $n$-cycle scenario with $d$ outcomes (Corollary 4.5). For fixed $(n,d)$, the number of $k$-order cycle distributions is given by
$\binom{d}{k}^{n}(k!)^{n-1}(k-1)!.$
Applying Stirling’s approximation to the case $k=d$ shows that the number of vertices grows super-exponentially in $d$, with leading-order behavior $\Theta((d/e)^{dn})$. This rapid growth, together with the general intractability of vertex enumeration for polyhedra [6], highlights that determining all extremal points of the simplicial polytope constitutes a computationally hard enumeration problem.
@@
Non-signaling distributions, and their generalization to simplicial distributions, can be utilized for computation. A typical example is the Mermin star contextuality scenario of [10], converted into a computation using the measurement-based quantum computation (MBQC) model [11]. The connection between contextuality and computational power is further analyzed in [12]. Our techniques are applicable to Bell scenarios that can be used for such computational purposes. In particular, the $(2,3,3)$ Bell scenario—two parties, each with three trichotomic measurements—can be collapsed to a measurement space obtained by gluing two cycle scenarios. Using this method, we are able to identify a contextual vertex (Section 5.3). Our main classification result for cycle scenarios with arbitrary outcomes opens up new applications to MBQC where the observables have an arbitrary number of outcomes. Such extensions have been studied in the literature, for example in [13].
More recently, Bell scenarios of the form (n,3,2)(n,3,2)—consisting of $n$ parties, each with three dichotomic measurements—have found applications in the classical simulation of quantum computation within the framework of quantum computation with magic states [14]. In this simulation model, the vertices of the polytope of simplicial distributions serve as classical keys, whose probabilistic mixtures represent the quantum state in the computation. Therefore, as a natural generalization to quantum measurements with $d$ outcomes [15], the enumeration of these vertices for Bell scenarios of the form $(n,d+1,d)$ seems to be crucial from the classical simulation perspective, providing a rigorous approach to identifying quantum advantage.
@@
Acknowledgments.
This work is supported by the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0002. The second and third authors acknowledge support from the Digital Horizon Europe project FoQaCiA, GA no. 101070558. The first and third authors are also supported by AFOSR FA9550-24-1-0257.
@@ 2.1 Simplicial distributions
  • An $n$-circle is a simplicial set $C^{(n)}$ specified by a sequence of pairwise distinct $1$-simplices $\sigma_{1},\cdots,\sigma_{n}$ satisfying $d_{0}(\sigma_{i})=d_{1}(\sigma_{i+1})$ for every $1\leq i\leq n-1$ and $d_{0}(\sigma_{n})=d_{1}(\sigma_{1})$. We will write $v_{i}=d_{1}(\sigma_{i})$.
@@
Figure 1: The 44-circle measurement space.
More generally, we will consider $1$-dimensional measurement spaces, i.e., basically directed graphs. In this case the simplicial set representing the measurement space is specified by its 0-simplices and $1$-simplices together with the face maps between them. We will write a distribution on an edge $\sigma$ as a $d\times d$-matrix $M$ where $M_{ab}=p^{ab}$. For convenience we index the rows and columns of such a matrix by $0,\cdots,d-1$. The marginals on the initial vertex and the terminal vertex are given by the sum of the rows and columns, respectively.
@@ 2.2 The vertex support
To prove (1) suppose that $\varphi\in\operatorname{supp}(p)$, i.e., $\delta^{\varphi}\preceq p$. By Corollary 2.6 this implies that $p=\delta^{\varphi}$, a contradiction since $p$ is contextual. Part (2) is proved in [17, Pro. 2.12]. ∎
@@ 2.3 Null-homotopy
We have a simplicial map $d_{0}:\text{Dec}^{0}(Y)\to Y$ that sends a simplex to its 0-th face. For $Y$ we will take the nerve space $N\mathbb{Z}_{d}$. The set of $n$-simplices of the nerve space consists of $n$-tuples of elements $(a_{1},a_{2},\cdots,a_{n})\in\mathbb{Z}_{d}^{n}$ together with the face and the degeneracy maps:
$d_{i}(a_{1},\cdots,a_{n})$ $=\left\{\begin{array}[]{ll}(a_{2},\cdots,a_{n})&i=0\\ (a_{1},\cdots,a_{i}+a_{i+1},\cdots,a_{n})&0<i<n\\ (a_{1},\cdots,a_{n-1})&i=n\end{array}\right$.
@@ 2.4 Vertices of faces
[…]
@@ 3.2 Morphisms of bundle scenarios
${\pi^{*}(E)}E{E^{\prime}}$${{X^{\prime}}}X{{X^{\prime}}}$$\scriptstyle{f}$$\scriptstyle{\alpha}$$\scriptstyle{g}$$\scriptstyle{\pi}$
@@
Part (2): Suppose that $p\in\operatorname{sDist}(f)$ is a vertex. If there exist $q,s\in\operatorname{sDist}(g)$ and $0<\alpha<1$ such that $i_{\ast}(p)=\alpha q+(1-\alpha)s$, then by Lemma 3.4 we have
@@
[…]
@@ 4 Cycle scenario with arbitrary outcome
According to Proposition 2.11 the map $\varphi$ is not null-homotopic. Let $p$ be a distribution in $\operatorname{Face}(\varphi)$, i.e., $D(\kappa)(p)=\delta^{\varphi}$. For $1\leq i\leq n-1$ we have $D(\kappa)(p)_{\sigma_{i}}=\delta^{0}$, i.e., $\sum_{a=0}^{k-1}p_{\sigma_{i}}^{aa}=1$, which means that $p_{\sigma_{i}}^{ab}=0$ if $b\neq a$. In addition, $D(\kappa)(p)_{\sigma_{n}}=\delta^{1}$, and thus $\sum_{a=0}^{k-1}p_{\sigma_{n}}^{a(a+1)}=1$, that means $p_{\sigma_{n}}^{ab}=0$ if $b\neq a+1$. So
@@
and $p_{\sigma_{n}}^{(k-1)0}=p_{d_{0}(\sigma_{n})}^{0}=p_{d_{1}(\sigma_{1})}^{0}=p_{\sigma_{1}}^{00}$. We conclude that $p_{\sigma_{i}}^{aa}=\frac{1}{k}$ for every $1\leq i\leq n-1$ and $p_{\sigma_{n}}^{a(a+1)}=\frac{1}{k}$. We proved that the only distribution in $\operatorname{Face}(\varphi)$ is the $k$-cycle distribution that corresponds to the following sequence
@@
With the explicit form of the $k$-order cycle distributions we can provide a formula for the number of vertices.
Corollary 4.5.
The number of vertices of the polytope of simplicial distributions on the $n$-cycle scenario with $d$-outcomes is given by the formula
$V_{n,d}=\sum_{k=1}^{d}\binom{d}{k}^{n}(k!)^{n-1}(k-1)!.$
Proof.
Let $V_{n,d,k}$ denote the number of $k$-order cycle distributions on the $n$-cycle scenario with $d$-outcomes. As described in Definition 4.1 a $k$-order distribution is specified by a $k\times n$ matrix over $\mathbb{Z}_{d}$:
$\begin{pmatrix}a_{1}^{(1)}&a_{2}^{(1)}&\cdots&a_{n}^{(1)}\\ a_{1}^{(2)}&a_{2}^{(2)}&\cdots&a_{n}^{(2)}\\ \vdots&\vdots&&\vdots\\ a_{1}^{(k)}&a_{2}^{(k)}&\cdots&a_{n}^{(k)}\\ \end{pmatrix}$
where in every column the entries are distinct. The number of such matrices is given by
$\binom{d}{k}^{n}(k!)^{n}.$
The first factor accounts for choosing $k$ distinct elements for each of the $n$ columns, and the second factor counts the number of ways to order the entries within each column. According to the formula for a $k$-order cycle distribution, each such matrix determines a unique distribution up to a cyclic permutation of its $k$ rows. Therefore, the number of distinct $k$-order cycle distributions is
$V_{n,d,k}=\frac{\binom{d}{k}^{n}(k!)^{n}}{k}=\binom{d}{k}^{n}(k!)^{n-1}(k-1)!.$
@@
Corollary 4.7 generalizes Theorem 1 from [9]. In that reference the authors describe the contextual (non-local) extremal distributions (vertices) on the 44-circle. Our result above extends this characterization to $n$-circle spaces with arbitrary outcomes.
@@ 5.1 Popescu-Rohrlich boxes
  • AA consists of a single edge ($1$-simplex) $\sigma_{1}$, and
@@
Figure 2: Decomposition of the 44-circle.
Since $A$ is a $1$-simplex and $B$ is a directed path obtained by gluing along 0-simplices it is known (see, e.g., [1, Example 3.11, Corollary 4.6]) that any distribution on $(A,\Delta_{\mathbb{Z}_{d}})$ and $(B,\Delta_{\mathbb{Z}_{d}})$ is noncontextual, thus the vertices are all deterministic. It follows that $\operatorname{Vsupp}(p|_{K}){\cong}\operatorname{supp}(p|_{K})$ for $K=A,B$. So one can see that
@@ 5.2 A scenario with trichotomic measurements
Figure 3: The measurement space decomposed into two parts given by 22-circles.
@@ 5.3 $(2,3,3)$ Bell scenario
$\Delta^{1}$$\Delta^{0}$$D(\Delta_{\mathbb{Z}_{3}})$$\scriptstyle{s^{0}}$$\scriptstyle{T}$$\scriptstyle{q}$
@@ 6 Conclusion
Our main contribution is to characterize all contextual extremal simplicial distributions of cycle scenarios with arbitrary outcomes by solving the vertex enumeration problem for the polytope of simplicial distributions on cycle scenarios with arbitrary outcomes. The simplicial distribution formulation aligns with the usual non-signaling formulation. Our result extends the known classification for binary outcomes to the case of arbitrary outcomes. Our approach uniquely benefits from novel homotopical methods, extending to scenarios formed by gluing cycle scenarios together. This gluing process is a crucial aspect of the theory of simplicial distributions when studying contextual distributions.
@@
Moreover, understanding contextual extremal distributions has applications to quantum advantage in the scheme of measurement-based quantum computation [11, 12, 24]. Our classification can be utilized to prove separation results in computational complexity, e.g., generalizing and extending results based on elevating a parity computer to a classically universal one. In addition, recent work has demonstrated that the vertices of polytopes of simplicial distributions arising from Bell scenarios play a central role in the classical simulation of quantum computation within the magic-state model [14]. Simplicial distributions provide the essential tools to investigate this kind of problem by providing otherwise impossible classification results, as we demonstrate in this paper. Our work opens the way for new results and brings a novel topological approach.